OpenStudy (brucebaner):

Amelia paid $252 for 28 cases of granola bars for the snack bar. The total cost varies directly with the number of cases purchased. How much will she pay for 10 cases of granola bars? $[_______]

OpenStudy (brucebaner):

@One098

OpenStudy (brucebaner):

is it 90?

OpenStudy (brucebaner):

@mathmath333

OpenStudy (brucebaner):

you there?

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{\normalsize \text{Amelia paid 252 for 28 cases of granola bars for the snack bar. } \hspace{.33em}\\~\\ \normalsize \text{The total cost varies directly with the number of cases purchased ,so } \hspace{.33em}\\~\\ \normalsize \text{let the total cost be x } \hspace{.33em}\\~\\ \normalsize \text{let the number of cases be y } \hspace{.33em}\\~\\ x\alpha y \hspace{.33em}\\~\\ x=ky \hspace{.33em}\\~\\ k=\dfrac{x}{y} \hspace{.33em}\\~\\ k=\dfrac{252}{28} =9 \hspace{.33em}\\~\\ x=9\cdot 10=90 \hspace{.33em}\\~\\ }\end{align}\)